# Large Language Models Meet Knowledge Graphs for Question Answering: Synthesis and Opportunities

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance on question-answering (QA) tasks because of their superior capabilities in natural language understanding and generation. However, LLM-based QA struggles with complex QA tasks due to poor reasoning capacity, outdated knowledge, and hallucinations. Several recent works synthesize LLMs and knowledge graphs (KGs) for QA to address the above challenges. In this survey, we propose a new structured taxonomy that categorizes the methodology of synthesizing LLMs and KGs for QA according to the categories of QA and the KG’s role when integrating with LLMs. We systematically survey state-of-the-art advances in synthesizing LLMs and KGs for QA and compare and analyze these approaches in terms of strength, limitations, and KG requirements. We then align the approaches with QA and discuss how these approaches address the main challenges of different complex QA. Finally, we summarize the advancements, evaluation metrics, and benchmark datasets and highlight open challenges and opportunities.
@@
Chuangtao Ma1, Yongrui Chen2, Tianxing Wu2, Arijit Khan1, Haofen Wang3 1Aalborg University, Denmark, 2Southeast University, China, 3Tongji University, China {chuma, arijitk}@cs.aau.dk, {yongruichen, tianxingwu}@seu.edu.cn, haofenwang@tongji.edu.cn
@@ 1 Introduction
Question answering (QA) plays a fundamental role in artificial intelligence, natural language processing, information retrieval, and data management areas since it has a wide range of applications such as text generation, chatbots, dialog generation, web search, entity linking, natural language query, fact-checking, etc. The pre-trained language models (PLMs) and recent LLMs have shown superior performance in several QA tasks such as KBQA (Knowledge bases QA), KGQA (Knowledge graph QA), CDQA (Closed domain QA), etc. However, PLM and LLM-based methods are incapable of handling complex QA scenarios due to the following limitations. (1) Limited complex reasoning capability: LLMs encapsulate very limited reasoning capability since they have been pre-trained with the tasks of predicting the next word in a text sequence. (2) Lack of up-to-date and domain-specific knowledge: LLMs are incapable of generating accurate and up-to-date responses for domain-specific QA because they have been pre-trained on the data and knowledge with a cutoff date. (3) Tendency to generate hallucinated content: LLMs usually generate hallucinated content due to the lack of factual verification and logical consistency checking.
Challenges. Retrieval augmented generation (RAG) (Mao et al., 2021) is proposed for open-domain QA by retrieving the relevant contexts from large documents, and several techniques, such as graph neural networks (GNNs) (Li et al., 2025b), have been investigated to enhance the retrieval coverage from passages. Although RAG-based QA can generate better responses in comparison to NoRAG-based QA, it still has a limited reasoning capacity and understanding of user interactions during complex QA. Complex QA usually involves knowledge interactions and fusion among data across modalities and sources, and an excellent understanding of complex queries and user interactions, whereas the RAG-based QA suffers a lot from the following technical challenges. (1) Knowledge conflicts: Conflicts occur due to the fusion of inconsistent and overlapping knowledge between LLMs and external sources in RAG-based QA that may further tend to generate inconsistent answers. (2) Poor relevance and quality of retrieved context: The accuracy of the generated answers in RAG-based QA largely depends on the relevance and quality of the retrieved context, where irrelevant context leads to incorrect results. (3) Lack of iterative and multi-hop reasoning: RAG-based QA struggles to generate accurate and explainable answers for questions requiring global and summarized contexts due to a lack of iterative.
@@
Taxonmy. We categorize the methodology of synthesizing LLMs and KGs and complex QA from different perspectives, and a structured taxonomy (details in Appendix § B) is given in Figure 1. Notably, these categories, divided from different perspectives, are non-exclusive, where a work might be classified into multiple categories from different perspectives. The online resources of this survey can be accessed via Github. [^1] https://github.com/machuangtao/LLM-KG4QA/
{forest}
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, font=, , [Large Language Models Meet Knowledge Graphs for QA, rotate=90, text width=6cm, fill=black!10 [Complex QA §2, text width=2.0cm, for tree=fill=cyan!20 [Multi-doc QA §2.1, text width=2.5cm [ KGP Wang et al. (2024d); CuriousLLM (Yang and Zhu, 2025); VisDom (Suri et al., 2024), text width=10.0cm, node options=align=left ] ] [Multi-modal QA §2.2, text width=2.5cm [ MMJG (Wang et al., 2022); RAMQA (Bai et al., 2025); KVQA (Dong et al., 2024b), text width=10.0cm, node options=align=left ] ] [Muti-hop QA §2.3, text width=2.5cm [ GraphLLM (Qiao et al., 2024); HOLMES (Panda et al., 2024); GMeLLo (Chen et al., 2024a), text width=10.0cm, node options=align=left ] ] [Muti-run and Conversational QA §2.4, text width=2.5cm [ CoRnNetA (Liu et al., 2024b); LLMs-Graph-Reasoning (Chen et al., 2024a); SELF-multi-RAG (Roy et al., 2024), text width=10.0cm, node options=align=left ] ] [Explainable QA §2.5, text width=2.5cm [ RoHT (Zhang et al., 2023); EXPLAIGNN (Christmann et al., 2023); RID (Feng et al., 2025), text width=10.0cm, node options=align=left ] ] [Temporal QA §2.6, text width=2.5cm [ TimeR4 (Qian et al., 2024); GenTKGQA (Gao et al., 2024); KG-IRAG (Yang et al., 2025), text width=10.0cm, node options=align=left ] ] ] [KGs as Background Knowledge §3.1, text width=2.0cm, for tree=fill=red!20 [Knowledge Integration and Fusion §3.1.1, text width=2.5cm [ InfuserKI (Wang et al., 2024a); KEFF (Zhao et al., 2025a); KnowLA (Luo et al., 2024b); KG-Adapter (Tian et al., 2024); GAIL (Zhang et al., 2024d), text width=10.0cm, node options=align=left ] ] [Retrieval Augmented Generation §3.1.2, text width=2.5cm [ PLRTQA (Alawwad et al., 2025); GNN-Ret (Li et al., 2025b); GRAG (Hu et al., 2024b); LEGO-GraphRAG (Cao et al., 2024); KG-RAG (Xu et al., 2024); KG2RAG (Zhu et al., 2025), text width=10.0cm, node options=align=left ] ] ] [KGs as Reasoning Guidelines §3.2, text width=2.0cm, for tree=fill=yellow!20 [Offline KG Guidelines §3.2.1, text width=2.5cm [ SR (Zhang et al., 2022); InfuserKI (Wang et al., 2024a); KnowLA (Luo et al., 2024b); KG-Adapter (Tian et al., 2024); GAIL (Zhang et al., 2024d); KELDaR (Li et al., 2024b), text width=10.0cm, node options=align=left ] ] [Online KG Guidelines §3.2.2, text width=2.5cm [ Oreo (Hu et al., 2022); KBIGER (Du et al., 2022); LLM-ARK (Huang, 2023); ToG (Sun et al., 2024a); ToG-2 (Ma et al., 2025b); KG-CoT (Zhao et al., 2024b), text width=10.0cm, node options=align=left ] ] [Agent-based KG Guidelines §3.2.3, text width=2.5cm [ KG-Agent (Jiang et al., 2024); ODA (Sun et al., 2024b); GREASELM (Zhang et al., 2021), text width=10.0cm, node options=align=left ] ] ] [KGs as Refiners and Validators §3.3, text width=2.0cm, for tree=fill=orange!20 [KG-Driven Filtering and Validation §3.3.1, text width=2.5cm [ ACT-Selection (Salnikov et al., 2023); Q-KGR (Zhang et al., 2024c); KG-Rank (Yang et al., 2024); KGR (Guan et al., 2024), text width=10.0cm, node options=align=left ] ] [KG-Augmented Output Refinement §3.3.2, text width=2.5cm [ EFSUM (Ko et al., 2024); InteractiveKBQA (Xiong et al., 2024); LPKG (Wang et al., 2024b), text width=10.0cm, node options=align=left ] ] ] [Advancements §3.4, text width=2.0cm, for tree=fill=blue!20 [Hybrid Method §3.4.1, text width=2.5cm, for tree=fill=blue!20 [ KG-RAG (Sanmartin, 2024); LongRAG (Zhao et al., 2024a); KG-Rank (Yang et al., 2024); FRAG (Zhao, 2024); KGQA (Ji et al., 2024), text width=10.0cm, node options=align=left ] ] [Optimization §3.4.2, text width=2.5cm, for tree=fill=blue!20 [ PG-RAG (Liang et al., 2024b); KGP (Wang et al., 2024d); SPOKE KG-RAG (Soman et al., 2024), text width=10.0cm, node options=align=left ] ] ] ]
@@
Contributions. Considering the popularity and mainstream adoption of both LLMs and KGs, and their wide applications in QA, our survey is timely. The contributions of this work are summarized below. (1) We introduce the structured taxonomy that categorizes state-of-the-art (SOTA) works on synthesizing LLMs+KGs for QA. (2) We review the SOTA works on synthesizing LLMs and KGs for QA in various categories and discuss the recent advanced topics in this field. (3) We create the alignments between various approaches in synthesizing LLMs and KGs and complex QA and highlight how these approaches address the specific challenges of different complex QA; (4) We discuss the current challenges and research opportunities on synthesizing LLMs and KGs for QA.
@@ 2 Complex QA
The methodology in synthesizing LLMs and KGs for QA has been exploited as follows.
@@ 2.3 Multi-hop QA
Multi-hop QA differs from simple QA, it usually involves multi-step reasoning to generate the final answers. The basic idea is to decompose the multi-hop questions into multiple single-hop questions, then generate the answers for each single-hop question, and finally integrate them (Linders and Tomczak, 2025). For instance, GraphLLM (Qiao et al., 2024) leverages LLMs to decompose the multi-hop question into several simple sub-questions and retrieve the sub-graphs via GNNs and LLMs to generate the answers for sub-questions based on graph reasoning. HOLME (Panda et al., 2024) utilizes a context-aware retrieved and pruned hyper-relational KG that is constructed based on the entity-document graph to enhance LLMs for generating the answers of multi-hop QA. To enable accurate fact retrieval and reasoning of LLMs for multi-hop QA, GMeLLo (Chen et al., 2024a) effectively integrates the explicit knowledge of KGs with the linguistic knowledge of LLMs by introducing the fact triple extraction, relation chain extraction, and query and answer generation.
2.4 Multi-run and Conversational QA
The challenge of multi-run and conversational QA lies in how to make the language model (LM) easily understand the questions and intermediate interactions. To make user interactions easily understood by machines, CoRnNetA (Liu et al., 2024b) introduces LLM-based question reformulations, reinforcement learning agents, and soft reward mechanisms to improve the interpretation of multi-turn interactions with KGs. The conversational QA involves several multi-run QA to refine and get accurate answers through multiple rounds of interactions. The knowledge aggregation module and graph reasoning are introduced for joint reasoning between the graph and LLMs (Jain and Lapata, 2024) to address the challenges of understanding the question and context for conversational QA. To improve the contextual understanding and the answer quality for conversational QA, SELF-multi-RAG (Roy et al., 2024) leverages LLMs to retrieve from the summarized conversational history and reuse the retrieved knowledge for augmentation.
@@ 2.5 Explainable QA
Explainable QA (XQA) aims to provide explanations for the generated answers based on the reasoning over the factual KGs. To effectively integrate the multiple sources of knowledge for XQA, RoHT (Zhang et al., 2023) introduces a two-stage XQA method that implements the probabilistic reasoning based on the constructed Hierarchical Question Decomposition Tree (HQDT) from the aggregated knowledge. To trace the provenance and improve the explainability of the answers, EXPLAIGNN (Christmann et al., 2023) constructs a heterogeneous graph from retrieved KB knowledge and user explanations and generates explanatory evidence based on GNN with question-level attention. RID (Feng et al., 2025) directly integrates the unsupervised retrieval with LLMs based on reinforcement learning-driven knowledge distillation.
2.6 Tempral QA
@@ 3 Approaches and Alignments
The strengths and limitations of the approach in synthesizing LLMs and KGs for QA are summarized (details in Table 2), and the research progress on the alignment of these approaches with the complex QA are discussed (details in Table 7).
@@ 3.1.1 Knowledge Integration and Fusion
As shown in Figure 2, the KGs and text are aligned via local subgraph extraction and entity linking, and then fed into the cross-model encoder to bidirectionally fuse text and KG to jointly train the language models for complex QA tasks (Yasunaga et al., 2022; Zhang et al., 2021). To address knowledge forgetting and noisy knowledge during knowledge integration, InfuserKI (Wang et al., 2024a) and KEFF (Zhao et al., 2025a) introduce the adaptive selection and knowledge enhancement filter, respectively, which selects the new knowledge and integrates it with LLMs. Fine-tuning LLMs with text and knowledge graphs can improve their performance of LLMs on specified tasks. For instance, KG-Adapter (Tian et al., 2024) improves parameter-efficient fine-tuning of LLMs by introducing a knowledge adaptation layer to LLMs. GAIL (Zhang et al., 2024d) fine-tunes LLMs for lightweight KGQA models based on retrieved SPARQL-question pairs from KGs.
@@ 3.1.2 Retrieval Augmented Generation
However, the mainstream RAG methods retrieve the relevant knowledge from the textual chunks, which ignores the structured information and inter-relations of these textual chunks. To mitigate this limitation, Graph RAG (Hu et al., 2024b) and KG-RAG (Sanmartin, 2024; Xu et al., 2024; Linders and Tomczak, 2025) are proposed. Instead of retrieving the knowledge from textual chunks, Graph RAG directly retrieves the relevant knowledge from graph data. GRAG (Hu et al., 2024b) retrieves the top-$k$ relevant subgraphs from the textual graph and then integrates the retrieved subgraphs with the query by aggregating and aligning the graph embeddings with text vectors based on GNNs. LEGO-GraphRAG (Cao et al., 2024) decomposes the retrieval process into subgraph-extraction, path-filtering, and path-refinement modules, thereby improving the reasoning capabilities of LLMs with retrieved knowledge. KG2RAG (Zhu et al., 2025) retrieves the relevant subgraph from KG and expanding the textual chunks with the retrieved KG for augmenting the generation.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance on question-answering (QA) tasks because of their superior capabilities in natural language understanding and generation. However, LLM-based QA struggles with complex QA tasks due to poor reasoning capacity, outdated knowledge, and hallucinations. Several recent works synthesize LLMs and knowledge graphs (KGs) for QA to address the above challenges. In this survey, we propose a new structured taxonomy that categorizes the methodology of synthesizing LLMs and KGs for QA according to the categories of QA and the KG’s role when integrating with LLMs. We systematically survey state-of-the-art methods in synthesizing LLMs and KGs for QA and compare and analyze these approaches in terms of strength, limitations, and KG requirements. We then align the approaches with QA and discuss how these approaches address the main challenges of different complex QA. Finally, we summarize the advancements, evaluation metrics, and benchmark datasets and highlight open challenges and opportunities.
@@
Chuangtao Ma1, Yongrui Chen2, Tianxing Wu2thanks: Corresponding authors., Arijit Khan3,  [^1] footnotemark: 1, Haofen Wang4 1Aalborg University, Denmark, 2Southeast University, China 3Bowling Green State University, USA, 4Tongji University, China chuma@cs.aau.dk, {yongruichen, tianxingwu}@seu.edu.cn, arijitk@bgsu.edu, haofenwang@tongji.edu.cn
@@ 1 Introduction
Question answering (QA) plays a fundamental role in artificial intelligence, natural language processing, information retrieval, and data management areas since it has a wide range of applications, such as text generation, chatbots, dialog generation, web search, entity linking, natural language query, fact-checking, etc. The pre-trained language models (PLMs) and recent LLMs have shown superior performance in several QA tasks such as KBQA (Knowledge bases QA), KGQA (Knowledge graph QA), CDQA (Closed domain QA), etc. However, PLM and LLM-based methods are incapable of handling complex QA scenarios due to the following limitations. (1) Limited complex reasoning capability: LLMs encapsulate very limited reasoning capability since they have been pre-trained with the tasks of predicting the next word in a text sequence. (2) Lack of up-to-date and domain-specific knowledge: LLMs are incapable of generating accurate and up-to-date responses for domain-specific QA because they have been pre-trained on the data and knowledge with a cutoff date. (3) Tendency to generate hallucinated content: LLMs usually generate hallucinated content due to the lack of factual verification and logical consistency checking.
Challenges. Retrieval augmented generation (RAG) (Mao et al., 2021) was proposed for open-domain QA by retrieving the relevant contexts from large documents, and several techniques, such as graph neural networks (GNNs) (Li et al., 2025b), have been investigated to enhance the retrieval coverage from passages. Although RAG-based QA can generate better responses in comparison to NoRAG-based QA, it still has limited capability for knowledge reasoning and understanding user interactions during complex QA. Complex QA usually involves knowledge interactions and fusion among data across modalities and sources, and an excellent understanding of complex queries and user interactions, whereas the RAG-based QA suffers a lot from the following technical challenges. (1) Knowledge conflicts: Conflicts occur due to the fusion of inconsistent and overlapping knowledge between LLMs and external sources in RAG-based QA that may further tend to generate inconsistent answers. (2) Poor relevance and quality of retrieved context: The accuracy of the generated answers in RAG-based QA largely depends on the relevance and quality of the retrieved context, where irrelevant context leads to incorrect results. (3) Lack of iterative and multi-hop reasoning: RAG-based QA struggles to generate accurate and explainable answers for questions requiring global and summarized contexts due to a lack of iterative.
@@
Taxonmy. We categorize the methodology of synthesizing LLMs and KGs and complex QA from different perspectives, and a structured taxonomy (details in Appendix § B) is given in Figure 1. The taxonomy from different perspectives aims to highlight the alignments between various LLM+KG approaches and different complex QA by discussing how the LLM+KG approaches with different roles of KG can address the challenges of complex QA. Notably, these categories, divided from different perspectives, are non-exclusive, where a work might be classified into multiple categories from different perspectives.
@@
Large Language Models Meet Knowledge Graphs for QA Complex QA §2 Multi-doc QA §2.1 KGP (Wang et al., 2024d); CuriousLLM (Yang and Zhu, 2025); VisDom (Suri et al., 2024) Multi-modal QA §2.2 MMJG (Wang et al., 2022); RAMQA (Bai et al., 2025); KVQA (Dong et al., 2024b) Muti-hop QA §2.3 GraphLLM (Qiao et al., 2024); HOLMES (Panda et al., 2024); GMeLLo (Chen et al., 2024b) Conversational QA §2.4 CoRnNetA (Liu et al., 2024b); LLMs-Graph-Reasoning (Chen et al., 2024b); SELF-multi-RAG (Roy et al., 2024) Explainable QA §2.5 RoHT (Zhang et al., 2023); EXPLAIGNN (Christmann et al., 2023); RID (Feng et al., 2025) Temporal QA §2.6 TimeR4 (Qian et al., 2024); GenTKGQA (Gao et al., 2024); KG-IRAG (Yang et al., 2025) KGs as Background Knowledge §3.1 Knowledge Integration and Fusion §3.1.1 InfuserKI (Wang et al., 2024a); KEFF (Zhao et al., 2025a); KnowLA (Luo et al., 2024b); KG-Adapter (Tian et al., 2024); GAIL (Zhang et al., 2024d) Retrieval Augmented Generation §3.1.2 PLRTQA (Alawwad et al., 2025); GNN-Ret (Li et al., 2025b); GRAG (Hu et al., 2024b); LEGO-GraphRAG (Cao et al., 2024); KG-RAG (Xu et al., 2024); KG2RAG (Zhu et al., 2025) KGs as Reasoning Guidelines §3.2 Offline KG Guidelines §3.2.1 SR (Zhang et al., 2022); InfuserKI (Wang et al., 2024a); KnowLA (Luo et al., 2024b); KG-Adapter (Tian et al., 2024); GAIL (Zhang et al., 2024d); KELDaR (Li et al., 2024b) Online KG Guidelines §3.2.2 Oreo (Hu et al., 2022); KBIGER (Du et al., 2022); LLM-ARK (Huang, 2023); ToG (Sun et al., 2024a; Ma et al., 2025b); KG-CoT (Zhao et al., 2024b); HippoRAG (Gutiérrez et al., 2024) Agent-based KG Guidelines §3.2.3 KG-Agent (Jiang et al., 2024); ODA (Sun et al., 2024b); GREASELM (Zhang et al., 2021); PoG (Chen et al., 2024a); ATOMR (Xin et al., 2025) KGs as Refiners and Validators §3.3 KG-Driven Filtering and Validation §3.3.1 ACT-Selection (Salnikov et al., 2023); Q-KGR (Zhang et al., 2024c); KG-Rank (Yang et al., 2024); KGR (Guan et al., 2024) KG-Augmented Output Refinement §3.3.2 EFSUM (Ko et al., 2024); InteractiveKBQA (Xiong et al., 2024); LPKG (Wang et al., 2024b) Advancements §3.4 Hybrid Method §3.4.1 KG-RAG (Sanmartin, 2024); LongRAG (Zhao et al., 2024a); KG-Rank (Yang et al., 2024); FRAG (Zhao, 2024); KGQA (Ji et al., 2024) Optimization §3.4.2 PG-RAG (Liang et al., 2024b); KGP (Wang et al., 2024d); SPOKE KG-RAG (Soman et al., 2024)
@@
Contributions. Considering the popularity and mainstream adoption of both LLMs and KGs, and their wide applications in QA, our survey is timely. The contributions of this work are summarized below. (1) We introduce the structured taxonomy that categorizes state-of-the-art (SOTA) works on synthesizing LLMs+KGs for QA. (2) We review the SOTA works on synthesizing LLMs and KGs for QA in various categories and discuss the recent advanced topics in this field. (3) We create the alignments between various approaches in synthesizing LLMs and KGs and complex QA, and highlight how these approaches address the specific challenges of different complex QA; (4) We discuss the current challenges and opportunities in synthesizing LLMs and KGs for QA. The online resources of this survey are available on Github [^1] https://github.com/machuangtao/LLM-KG4QA.
@@ 2 Complex QA
Complex QA usually involves question decomposition, knowledge fusion among data across modalities and sources, where complex knowledge reasoning is required to generate accurate answers. The methodology in synthesizing LLMs and KGs for complex QA has been exploited as follows.
@@ 2.3 Multi-hop QA
Multi-hop QA differs from simple QA, which usually involves multi-step reasoning to generate the final answers. The basic idea is to decompose the multi-hop questions into multiple single-hop questions, then generate the answers for each single-hop question, and finally integrate them (Linders and Tomczak, 2025). For instance, GraphLLM (Qiao et al., 2024) leverages LLMs to decompose the multi-hop question into several simple sub-questions and retrieve the sub-graphs via GNNs and LLMs to generate the answers for sub-questions based on graph reasoning. HOLME (Panda et al., 2024) utilizes a context-aware retrieved and pruned hyper-relational KG that is constructed based on the entity-document graph to enhance LLMs for generating the answers of multi-hop QA. To enable accurate factual knowledge retrieval and reasoning of LLMs for multi-hop QA, GMeLLo (Chen et al., 2024b) effectively integrates the explicit knowledge of KGs with the linguistic knowledge of LLMs by introducing the fact triple extraction, relation chain extraction, and query and answer generation.
2.4 Conversational QA
The challenge of conversational QA lies in how to make the language model (LM) easily understand the questions and intermediate interactions. To make user interactions easily understood by machines, CoRnNetA (Liu et al., 2024b) introduces an LLM-based question reformulation, reinforcement learning agent, and soft reward mechanism to improve the interpretation of multi-turn interactions with KGs. The conversational QA involves several multi-run QA to refine and get accurate answers through multiple rounds of interactions. The knowledge aggregation module and graph reasoning are introduced for joint reasoning between the graph and LLMs (Jain and Lapata, 2024) to address the challenges of understanding the question and context for conversational QA. To improve the contextual understanding and the answer quality for conversational QA, SELF-multi-RAG (Roy et al., 2024) leverages LLMs to retrieve from the summarized conversational history and reuse the retrieved knowledge for augmentation.
@@ 2.5 Explainable QA
Explainable QA (XQA) aims to provide explanations for the generated answers based on the reasoning over the factual KGs. To effectively integrate the multiple sources of knowledge for XQA, RoHT (Zhang et al., 2023) introduces a two-stage XQA method that implements the probabilistic reasoning based on the constructed Hierarchical Question Decomposition Tree (HQDT) from the aggregated knowledge. To trace the provenance and improve the explainability of the answers, EXPLAIGNN (Christmann et al., 2023) constructs a heterogeneous graph from retrieved knowledge and user explanations and generates explanatory evidence based on GNN with question-level attention. RID (Feng et al., 2025) directly integrates the unsupervised retrieval with LLMs based on reinforcement learning-driven knowledge distillation.
2.6 Temporal QA
@@ 3 Approaches and Alignments
The strengths and limitations of the approach in synthesizing LLMs and KGs for QA are summarized (details in Table 2), and the research progress on the alignment of these approaches with the complex QA is discussed (details in Table 7).
@@ 3.1.1 Knowledge Integration and Fusion
As shown in Figure 2, the KGs and text are aligned via local subgraph extraction and entity linking, and then fed into the cross-model encoder to bidirectionally fuse text and KG to jointly train the language models for complex QA tasks (Yasunaga et al., 2022; Zhang et al., 2021). To address knowledge forgetting and noisy knowledge during knowledge integration, InfuserKI (Wang et al., 2024a) and KEFF (Zhao et al., 2025a) introduce the adaptive selection and knowledge enhancement filter, respectively, which selects the new knowledge and integrates it with LLMs. Fine-tuning LLMs with text and knowledge graphs can improve their performance on specified tasks. For instance, KG-Adapter (Tian et al., 2024) improves parameter-efficient fine-tuning of LLMs by introducing a knowledge adaptation layer to LLMs. GAIL (Zhang et al., 2024d) fine-tunes LLMs for lightweight KGQA models based on retrieved SPARQL-question pairs from KGs.
@@ 3.1.2 Retrieval Augmented Generation
However, the mainstream RAG methods retrieve the relevant knowledge from the textual chunks, which ignore the structured information and inter-relations of these textual chunks. To mitigate this limitation, Graph RAG (Hu et al., 2024b) and KG-RAG (Sanmartin, 2024; Xu et al., 2024; Linders and Tomczak, 2025) are proposed. Instead of retrieving the knowledge from textual chunks, Graph RAG directly retrieves the relevant knowledge from graph data. GRAG (Hu et al., 2024b) retrieves the top-$k$ relevant subgraphs from the textual graph and then integrates the retrieved subgraphs with the query by aggregating and aligning the graph embeddings with text embeddings based on GNNs. LEGO-GraphRAG (Cao et al., 2024) decomposes the retrieval process into subgraph-extraction, path-filtering, and path-refinement modules, thereby improving the reasoning capabilities of LLMs with retrieved knowledge. KG2RAG (Zhu et al., 2025) retrieves the relevant subgraph from KG and expanding the textual chunks with the retrieved KG for augmenting the generation.